# A Novel Deep Hybrid Framework with Ensemble-Based Feature Optimization for Robust Real-Time Human Activity Recognition

### v1

@@ Abstract
Human Activity Recognition (HAR) plays a pivotal role in various applications, including smart surveillance, healthcare, assistive technologies, sports analytics, etc. However, HAR systems still face critical challenges, including high computational costs, redundant features, and limited scalability in real-time scenarios. An optimized hybrid deep learning framework is introduced that integrates a customized InceptionV3, an LSTM architecture, and a novel ensemble-based feature selection strategy. The proposed framework first extracts spatial descriptors using the customized InceptionV3 model, which captures multilevel contextual patterns, region homogeneity, and fine-grained localization cues. The temporal dependencies across frames are then modeled using LSTMs to effectively encode motion dynamics. Finally, an ensemble-based genetic algorithm with Adaptive Dynamic Fitness Sharing and Attention (ADFSA) is employed to select a compact and optimized feature set by dynamically balancing objectives such as accuracy, redundancy, uniqueness, and complexity reduction. Consequently, the selected feature subsets, which are both diverse and discriminative, enable various lightweight machine learning classifiers to achieve accurate and robust HAR in heterogeneous environments. Experimental results on the robust UCF-YouTube dataset, which presents challenges such as occlusion, cluttered backgrounds, motion dynamics, and poor illumination, demonstrate good performance. The proposed approach achieves 99.65% recognition accuracy, reduces features to as few as 7, and enhances inference time. The lightweight and scalable nature of the HAR system supports real-time deployment on edge devices such as Raspberry Pi, enabling practical applications in intelligent, resource-aware environments, including public safety, assistive technology, and autonomous monitoring systems.
@@ 1 Introduction
Human activity recognition (HAR) plays a pivotal role in the development of smart cities, particularly in the context of video-based surveillance. Accurate identification of human behavior can help law enforcement and public safety agencies detect irregular activities, such as accidents, property damage, or illicit behaviors such as fighting and theft. Early research in HAR focused primarily on controlled environments with a single actor, but recent studies have shifted toward more realistic and challenging scenarios. These include uncontrolled video datasets with complex backgrounds, varying illumination, camera movement, and occlusion, making accurate recognition considerably more difficult [1]. Improving pedestrian detection and tracking in such dynamic surveillance situations remains a significant challenge.
@@
Although Convolutional Neural Networks (CNNs) have demonstrated remarkable performance in image classification and object detection, they are inherently limited to processing individual frames. They cannot directly capture temporal dynamics in video sequences [2]. To address this, two-stream CNNs were proposed, in which appearance and motion information (e.g., optical flow) are processed separately. Although they achieve good recognition performance, they incur high computational costs, limiting their suitability for large-scale or real-time applications [3, 4]. More recently, 3D CNNs have been introduced to directly extract spatiotemporal information from raw video streams [5, 6]. To enhance temporal modeling, recurrent architectures such as RNNs and LSTMs have also been incorporated [7, 8]. These approaches suffer from high computational cost, especially when processing long video sequences, which limits their applicability for large-scale and real-time scenarios.
@@
To address these limitations, researchers have explored feature fusion and selection strategies for HAR. Examples include combining deep spatial and motion representations using spatial–temporal attention mechanisms [9] and combining handcrafted multiview features with feature selection based on entropy-variance [10]. Although these methods report promising performance, they often rely on hand-crafted components and produce high-dimensional feature representations, leading to increased computational complexity and limited scalability. Efficient selection of discriminative features is therefore essential to improve inference speed without compromising accuracy.
@@
In this work, we propose a novel hybrid HAR framework that integrates a customized InceptionV3 CNN for spatial feature extraction with an LSTM network for temporal modeling. The CNN captures discriminative spatial patterns from individual frames, while the LSTM models temporal dependencies across sequences. Designed for real-time video processing, the framework handles challenges such as occlusion, cluttered backgrounds, motion blur, and multiview variations, as exemplified by the UCF YouTube action dataset [11]. Both models are end-to-end trainable and optimized via transfer learning and fine-tuning to ensure efficient inference. Furthermore, to reduce feature redundancy and computational overhead, a genetic algorithm (GA)–based feature selection strategy is applied, selecting only the most discriminative spatiotemporal features for final classification. By eliminating the need for costly preprocessing steps such as optical flow or skeleton extraction, the proposed approach achieves high accuracy while maintaining practical real-time applicability.
@@
  • This study proposes a hybrid DL framework that synergistically integrates spatial customized InceptionV3 and temporal (LSTM) modeling for HAR from video data. The customized InceptionV3 integrates region-based (average pooling) and boundary-aware (max pooling) operations after each convolutional block to enhance region homogeneity, suppress noise, and capture discriminative local features, while improving robustness through down-sampling. Moreover, the enhanced InceptionV3 extracts fine-grained localization, contextual cues, and human positioning within frames, which a customized LSTM sequentially processes to model inter-frame temporal dependencies and motion dynamics with high fidelity.
  • A novel ensemble-based GA-driven feature selection strategy, ADFSA, is employed to achieve compact and discriminative feature subsets. The ADFSA mechanism dynamically reweights multiple objectives: classification accuracy, redundancy minimization, uniqueness enhancement, and complexity reduction through attention-guided adaptation, thereby accelerating convergence and mitigating local stagnation during the evolutionary process.
@@ 2 Related Work
Recent studies show a growing interest in leveraging advanced deep learning architectures for human action recognition, including CNN-based models, transfer learning approaches, and, more recently, Transformer-based sequence learning. For example, the BT-LSTM model proposed in [12] achieved an accuracy of 85.3% on the UCF11 dataset. Considerable progress has since been made with various deep learning approaches that attempt to jointly capture spatial and temporal information. Early methods such as the CNN–BiLSTM network [13] and the deep autoencoder combined with CNN [14] reported recognition accuracies of 92.84% and 96.2%, respectively. However, these end-to-end architectures require large-scale training datasets and entail high computational costs.
To improve spatial representation, subsequent studies explored handcrafted or hybrid descriptors. For instance, the key-frame driven KFDI approach [15] achieved only 79.4% due to limited spatial discrimination. Although more advanced techniques, such as the dilated CNN–BiLSTM with residual blocks [16] and the local–global feature fusion method with QSVM [17], achieved 89.01% and 82.6% accuracy, respectively, their generalization capability remained limited when applied to more complex and unconstrained activities.
More recently, fully 3D networks and transformer-based encoders have been introduced to capture long-term temporal dependencies. The 3D‐CNN in [18] achieved 85.20%, and the ConvNeXt–TCN model in [19] reached 97.73%, demonstrating the importance of temporal aggregation. Similarly, hybrid deep models such as VGG–BiGRU [20] and ViT‐ReT [21] achieved 93.38% and 92.4%, respectively. Despite these advances, most of these approaches employ complex, parameter-heavy architectures and retain high-dimensional feature representations, which leads to increased inference time and reduced efficiency.
Although the bidirectional LSTM model presented in [22] achieved a high recognition accuracy of 99.2%, it still relies on full feature vectors without any dimensionality reduction, limiting its suitability for real-time or edge deployment. In contrast, the proposed method attains a higher accuracy of 99.58% while substantially reducing the feature dimensionality. By extracting spatial features using a lightweight CNN and subsequently applying efficient machine learning classifiers to a compact set of selected features, the proposed framework provides a robust and computationally efficient solution that outperforms existing methods in both accuracy and inference efficiency. A summary of reported previous work is presented in Table 1.
@@
To overcome these shortcomings, this study proposes an ensemble ADFSA approach. The method integrates multiple optimization objectives, including accuracy, redundancy, uniqueness, and complexity, that dynamically adjust their relative importance throughout the evolutionary process. Additionally, by leveraging a two-stage deep feature extraction pipeline (customized InceptionV3 + LSTM), the method effectively captures both spatial and temporal representations. This hybrid design enables the proposed approach to achieve better recognition performance, faster inference, and reduced model complexity, making it well-suited for real-time HAR applications.
@@ 3 Methodology
This section introduces the proposed novel hybrid deep feature extraction and selection methodology for HAR. In video-based HAR systems, human activities are mainly recognized from continuous video streams acquired through surveillance cameras or other visual sensors. The overall proposed framework for hybrid deep feature extraction and selection is shown in Figure 1, and the detailed explanation and presentation are provided in subsequent sections.
Figure 1: Overall Block Diagram of Proposed Methodology.
3.1 Frames Extraction from Video Clip
@@ 3.2 Pre-processing
3.3 Data Augmentation
To enhance the diversity and generalization of the model, data augmentation techniques are applied during the pre-processing phase. These augmentations simulate real-world variations and prevent overfitting by introducing randomness into the training data. Common transformations include random horizontal flipping, slight rotations, zooming, and brightness adjustments, all applied within controlled bounds to preserve the semantic content of human activities. These augmentations are executed on a per-frame basis before the frames are fed into the feature extractor. By increasing variability in spatial patterns without altering temporal coherence, data augmentation enhances the robustness and reliability of temporal feature learning in downstream LSTM-based modeling.
3.4 Deep Hybrid Convolutional Neural Network
This paper proposes a deep hybrid architecture that combines a customized InceptionV3 model with an LSTM network for effective human activity recognition. The inception-V3 model is employed to extract rich spatial features from individual video frames, while the LSTM network captures the temporal dependencies across frame sequences. This integration enables the model to learn both appearance-based and motion-based patterns, making it well-suited for complex video-based activity classification tasks.
3.4.1 Customized Inception-V3 Network
The proposed model utilizes a customized inception-V3 architecture, specifically tailored to enhance spatial feature extraction for human activity recognition tasks. Modifications to the standard Inception modules include the strategic integration of pooling operations within individual branches to improve feature diversity and hierarchical abstraction. In the first Inception block, a Max-pooling (3×3) operation is added alongside Convolutional branches to capture dominant spatial activations early in the network. The second block is enhanced with both Average-pooling (3×3) after three Convolutional layers and Max-pooling (3×3) after two Convolutional layers, promoting multi-scale feature fusion. The third and fourth blocks continue this pattern with the pooling operations placed at various depths within the branches to extract increasingly abstract representations. The network concatenates all the blocks and finally passes through an 8×8 Final Global Average Pooling layer, reducing the feature map to a compact 1×1×1024 tensor. This refined design facilitates the generation of a robust 1024-dimensional feature vector that encodes discriminative spatial information crucial for downstream temporal modeling.
@@ 3.4.1 Customized Inception-V3 Network
In parallel, max pooling is performed (Equation 4) to capture the most salient features within the same region. It learns region-specific variations and their characteristic boundaries [23]. Max-pooling selects the strongest activation functions within the pooling region, emphasizing dominant, discriminative features by capturing key motion cues such as joint positions or sharp limb movements. It increases translation invariance to small spatial shifts and reduces irrelevant background by discarding weaker activations [24]. Additionally, it speeds up computation by shrinking feature maps without losing essential details, ensuring that critical actions like the leg motion in a “soccer kick” remain prominent. The outputs from multiple parallel branches (such as different filter sizes or pooling strategies) are then concatenated depth-wise to form a unified feature map $Y$, as shown in Equation 5, where the total number of channels $D$ is the sum of the individual branch depths. This concatenated feature map is flattened into a 1D vector $v$ (Equation 7) to prepare it for fully connected layers. The first fully connected layer transforms $v$ into a higher-level abstract representation $h_{k}$ using learned weights and a non-linear activation function $f$, as illustrated in Equation 6. This is followed by a second dense layer (Equation 8), which refines these features into final activations $f_{n}$. The complete 1024-dimensional feature vector $\mathbf{f}$, shown in Equation 9, serves as the final spatial representation extracted from the image and is subsequently passed to the temporal modeling module, such as an LSTM network.
3.4.2 Spatio-Temporal Representation via Customized LSTM
Following the extraction of spatial features from each frame using the customized InceptionV3 model, sequences of 20 consecutive 1024-dimensional vectors, representing one video segment, are used to train an LSTM network. The LSTM is designed with 128 hidden units and is trained to model temporal dependencies across the sequential spatial features, learning dynamic transitions and contextual patterns associated with various human activities. After training, the LSTM is detached from its classification objective and repurposed as a temporal feature extractor. For each input sequence of shape $(20,1024)$, the trained LSTM processes the temporal structure and outputs a 128-dimensional feature vector corresponding to its final hidden state. This vector serves as a compact spatial–temporal representation, enriched with both appearance-based and motion-based information, and is used in subsequent stages such as feature fusion or classification.

### v2

@@ Abstract
Human Activity Recognition (HAR) plays a pivotal role in various applications, including smart surveillance, healthcare, assistive technologies, sports analytics, etc. However, HAR systems still face critical challenges, including high computational costs, redundant features, and limited scalability in real-time scenarios. An optimized hybrid deep learning framework is introduced that integrates a customized InceptionV3, an Attention-Augmented AA-LSTM architecture, and a novel ensemble-based feature selection strategy. The proposed framework first extracts spatial descriptors using the customized InceptionV3 model, which captures multilevel contextual patterns, region homogeneity, and fine-grained localization cues. The temporal dependencies across frames are then modeled using AA-LSTMs to effectively encode motion dynamics. Finally, an ensemble-based genetic algorithm with Adaptive Dynamic Fitness Sharing and Attention (ADFSA) is employed to select a compact and optimized feature set by dynamically balancing objectives such as accuracy, redundancy, uniqueness, and complexity reduction. Consequently, the selected feature subsets, which are both diverse and discriminative, enable various lightweight machine learning classifiers to achieve accurate and robust HAR in heterogeneous environments. Experimental results on the robust UCF-YouTube dataset, which presents challenges such as occlusion, cluttered backgrounds, motion dynamics, and poor illumination, demonstrate good performance. The proposed framework achieves 99.65% recognition accuracy, reduces features to as few as 7, and enhances inference time. The lightweight and scalable nature of the HAR system supports real-time deployment on edge devices such as Raspberry Pi, enabling practical applications in intelligent, resource-aware environments, including public safety, assistive technology, and autonomous monitoring systems.
@@ 1 Introduction
Human activity recognition (HAR) plays a pivotal role in the development of smart cities, particularly in the context of video-based surveillance. Accurate identification of human behavior can help law enforcement and public safety agencies detect irregular activities ranging from accidents and property damage to illicit behaviors such as fighting and theft. Early research in HAR focused primarily on controlled environments with a single actor, but recent studies have shifted toward more realistic and challenging scenarios. These include uncontrolled video datasets with complex backgrounds, varying illumination, camera movement, and occlusion, making accurate recognition considerably more difficult [1]. Improving pedestrian detection and tracking in such dynamic surveillance situations remains a significant challenge.
@@
Although Convolutional Neural Networks (CNNs) have demonstrated remarkable performance in image classification and object detection, they are inherently limited to processing individual frames. They cannot directly capture temporal dynamics in video sequences [2]. To address this, two-stream CNNs were presented, in which appearance and motion information (e.g., optical flow) are processed separately. Although they achieve good recognition performance, they incur high computational costs, limiting their suitability for large-scale or real-time applications [3, 4]. More recently, 3D CNNs have been introduced to directly extract spatiotemporal information from raw video streams [5, 6]. To enhance temporal modeling, recurrent architectures such as RNNs and LSTMs have also been incorporated [7, 8]. These approaches suffer from high computational cost, especially when processing long video sequences, which limits their applicability for large-scale and real-time scenarios.
Recently, Transformer-based models have become popular in human activity recognition. For video data, researchers have combined 3D convolution with attention-based Transformers to better capture both spatial and temporal patterns, while efficiency-focused architectures have enabled lightweight, real-time processing in open-set HAR tasks [9, 10]. In the wearable sensor domain, methods like MoPFormer, SETransformer, and P2LHAP have introduced novel innovations, including tokenization into motion primitives, channel-wise attention, and sequence forecasting, achieving improved accuracy, interpretability, and generalization across datasets [11, 12]. Lightweight, end-to-end Transformers for smartphone IMU data also match or exceed CNN–LSTM baselines with fewer parameters and strong performance [13]. Despite these gains, recent surveys note that Transformer models often entail heavy computation, dependency on large datasets, and limitations when deployed on resource-constrained devices, posing challenges to their broader adoption [14].
@@
To address these limitations, researchers have explored feature fusion and selection strategies for HAR. Promising performance achieved by incorporating deep spatial and motion representations using spatial–temporal attention mechanisms [15] and combining hand-crafted multi-view features with feature selection (FS) based on entropy variability [16]. They often rely on hand-crafted components and produce high-dimensional feature representations, leading to increased computational complexity and limited scalability. Efficient selection of discriminative features is therefore essential to improve inference speed without compromising accuracy.
@@
In this work, we propose a novel hybrid HAR framework that integrates a customized InceptionV3 CNN for spatial feature extraction with an Attention-Augmented (AA-LSTM) network for temporal modeling. The CNN captures discriminative spatial patterns from individual frames, while the AA-LSTM models temporal dependencies across sequences. Designed for real-time video processing, the framework handles challenges such as occlusion, cluttered backgrounds, motion blur, and multiview variations, as exemplified by the UCF YouTube action dataset [17]. Both models are end-to-end trainable and optimized via transfer learning and fine-tuning to ensure efficient inference. Furthermore, to reduce feature redundancy and computational overhead, a genetic algorithm (GA)–based FS strategy is applied, selecting only the most discriminative spatiotemporal features for final classification. By eliminating the need for costly preprocessing steps such as optical flow or skeleton extraction, the proposed framework achieves high accuracy while maintaining practical real-time applicability.
@@
  • This study proposes a hybrid DL framework that synergistically integrates spatial customized InceptionV3 and temporal AA-LSTM modeling for HAR from video data. The customized InceptionV3 integrates region-based (average pooling) and boundary-aware (max pooling) operations after each convolutional block to enhance region homogeneity, suppress noise, and capture discriminative local features, while improving robustness through down-sampling. Moreover, the enhanced InceptionV3 extracts fine-grained localization, contextual cues, and human positioning within frames, which an AA-LSTM sequentially processes to model inter-frame temporal dependencies and motion dynamics with high fidelity.
  • A novel ensemble-based GA-driven FS strategy, ADFSA, is employed to achieve compact and discriminative feature subsets. The ADFSA mechanism dynamically reweights multiple objectives: classification accuracy, redundancy minimization, uniqueness enhancement, and complexity reduction through attention-guided adaptation, thereby accelerating convergence and mitigating local stagnation during the evolutionary process.
@@ 2 Related Work
Recent research has increasingly emphasized advanced DL architectures for HAR, spanning CNN-based frameworks, transfer learning paradigms, and, more recently, Transformer-driven sequence modeling. The BT-LSTM model [18], for instance, achieved 85.3% accuracy on the UCF11 dataset, highlighting the importance of temporal modeling but exhibiting limited generalization in unconstrained scenarios. To jointly capture spatial and temporal dependencies, several end-to-end architectures have been explored. The CNN–BiLSTM framework [19] attained 92.84% accuracy, while the deep autoencoder–CNN approach [20] reached 96.2%. However, these models required large-scale training data and involved substantial computational costs, reducing their feasibility in real-time applications.
Efforts to strengthen spatial representation also included handcrafted and hybrid descriptors. The key-frame driven KFDI method [21] achieved only 79.4%, largely due to weak spatial discrimination. More sophisticated designs, such as the dilated CNN–BiLSTM with residual blocks [22] and the local–global feature fusion with QSVM [23], achieved 89.01% and 82.6%, respectively. However, these methods exhibited reduced robustness when applied to complex activities with occlusions, cluttered backgrounds, and motion variations.
More recently, fully 3D CNN and Transformer-based encoders have been introduced to address long-term temporal modeling. The 3D-CNN [24] reported 85.20% accuracy, whereas the ConvNeXt–TCN framework [25] achieved 97.73%, underscoring the role of temporal aggregation. Hybrid architectures such as VGG–BiGRU [26] and ViT-ReT [27] attained 93.38% and 92.4% accuracy, respectively. However, despite these advancements, most of these models employed parameter-heavy architectures and retained high-dimensional feature representations, resulting in increased inference latency and limited applicability on resource-constrained platforms.
Additionally, the bidirectional LSTM model [28] reported a high recognition accuracy of 99.2%, yet it relied on full feature vectors without any dimensionality reduction. This dependence on high-dimensional features limited its scalability and efficiency in real-time or edge-deployment environments. Overall, prior studies have demonstrated significant progress but continue to face challenges related to computational efficiency, redundancy in feature representation, and deployment feasibility. A concise summary of existing approaches is presented in Table 1.
@@
Despite notable progress, existing HAR models still face challenges related to computational efficiency, redundancy in feature representation, and deployment feasibility on real-time or edge platforms. This highlights the need for lightweight yet effective frameworks that can balance recognition accuracy with scalability and efficiency. To overcome these shortcomings, this study proposes a novel deep hybrid framework with ensemble ADFSA FS scheme. The method integrates multiple optimization objectives, including accuracy, redundancy, uniqueness, and complexity, that dynamically adjust their relative importance throughout the evolutionary process. Additionally, by leveraging a two-stage deep feature extraction pipeline (customized InceptionV3 + AA-LSTM), the method effectively captures both spatial and temporal representations. This hybrid design enables the proposed framework to achieve better recognition performance, faster inference, and reduced model complexity, making it well-suited for real-time HAR applications.
@@ 3 Methodology
This study presents a novel hybrid deep feature extraction and selection framework for HAR, structured into four sequential phases:
  1. 1.
    Data Preprocessing: Video data is decomposed into frames, uniformly sampled, resized to $224\times 224$ pixels, normalized, and augmented through controlled transformations to enhance robustness and mitigate overfitting.
  2. 2.
    Spatial Feature Extraction: A customized InceptionV3 captures fine-grained spatial descriptors, including contextual cues, region homogeneity, and discriminative local patterns.
  3. 3.
    Temporal Modeling: An AA-LSTM encodes inter-frame dependencies and motion dynamics with high temporal fidelity.
  4. 4.
    Feature Selection and Classification: An ensemble-based genetic algorithm with ADFSA identifies compact, non-redundant, and highly discriminative feature subsets, which are subsequently classified using lightweight machine learning models.
This phased pipeline shown in Figure 1, ensures a balance between accuracy, compactness, and computational efficiency, supporting real-time deployment on edge devices. The framework is further benchmarked against state-of-the-art methods to validate its effectiveness.
Figure 1: Overall Block Diagram of Proposed Framework.
3.1 Data Preprocessing
3.1.1 Frames Extraction from Video Clip
@@ 3.1.2 Image resizing
3.1.3 Data Augmentation
To enhance the diversity and generalization of the model, data augmentation techniques are applied during the preprocessing phase. These augmentations simulate real-world variations and prevent overfitting by introducing randomness into the training data [29]. Common transformations include random horizontal flipping, slight rotations, zooming, and brightness adjustments, all applied within controlled bounds to preserve the semantic content of human activities. These augmentations are executed on a per-frame basis before the frames are fed into the feature extractor. By increasing variability in spatial patterns without altering temporal coherence, data augmentation enhances the robustness and reliability of temporal feature learning in downstream AA-LSTM-based modeling.
3.2 Feature Extraction via Deep Hybrid Convolutional Neural Network
This paper proposes a deep hybrid architecture that combines a customized InceptionV3 model with an AA-LSTM network for effective human activity recognition. The inception-V3 model is employed to extract rich spatial features from individual video frames, while the AA-LSTM network captures the temporal dependencies across frame sequences. This integration enables the model to learn both appearance-based and motion-based patterns, making it well-suited for complex video-based activity classification tasks.
3.2.1 Spatial feature extraction via Customized Inception-V3 Network
The proposed framework utilizes a customized inception-V3 architecture, specifically tailored to enhance spatial feature extraction for human activity recognition tasks. Modifications to the standard Inception modules include the strategic integration of pooling operations within individual branches to improve feature diversity and hierarchical abstraction. In the first Inception block, a Max-pooling (3×3) operation is added alongside Convolutional branches to capture dominant spatial activations early in the network. The second block is enhanced with both Average-pooling (3×3) after three Convolutional layers and Max-pooling (3×3) after two Convolutional layers, promoting multi-scale feature fusion. The third and fourth blocks continue this pattern with the pooling operations placed at various depths within the branches to extract increasingly abstract representations. The network concatenates all the blocks and finally passes through an 8×8 Final Global Average Pooling layer, reducing the feature map to a compact 1×1×1024 tensor. This refined design facilitates the generation of a robust 1024-dimensional feature vector that encodes discriminative spatial information crucial for downstream temporal modeling.
@@ 3.2.1 Spatial feature extraction via Customized Inception-V3 Network
In parallel, max pooling is performed (Equation 4) to capture the most salient features within the same region. It learns region-specific variations and their characteristic boundaries [30, 31]. Max-pooling selects the strongest activation functions within the pooling region, emphasizing dominant, discriminative features by capturing key motion cues such as joint positions or sharp limb movements. It increases translation invariance to small spatial shifts and reduces irrelevant background by discarding weaker activations [32]. Additionally, it speeds up computation by shrinking feature maps without losing essential details, ensuring that critical actions like the leg motion in a “soccer kick” remain prominent. The outputs from multiple parallel branches (such as different filter sizes or pooling strategies) are then concatenated depth-wise to form a unified feature map $Y$, as shown in Equation 5, where the total number of channels $D$ is the sum of the individual branch depths. This concatenated feature map is flattened into a 1D vector $v$ (Equation 7) to prepare it for fully connected layers. The first fully connected layer transforms $v$ into a higher-level abstract representation $h_{k}$ using learned weights and a non-linear activation function $f$, as illustrated in Equation 6. This is followed by a second dense layer (Equation 8), which refines these features into final activations $f_{n}$. The complete 1024-dimensional feature vector $\mathbf{f}$, shown in Equation 9, serves as the final spatial representation extracted from the image and is subsequently passed to the temporal modeling module, such as an AA-LSTM network.
3.2.2 Spatio-Temporal Representation via Attention-Augmented LSTM
Following the extraction of spatial features from each frame using the customized InceptionV3 model, sequences of 20 consecutive 1024-dimensional vectors, representing one video segment, are used to train an AA-LSTM network. The AA-LSTM is designed with 128 hidden units and is trained to model temporal dependencies across the sequential spatial features, learning dynamic transitions and contextual patterns associated with various human activities. After training, the AA-LSTM is detached from its classification objective and repurposed as a temporal feature extractor. For each input sequence of shape $(20,1024)$, the trained AA-LSTM processes the temporal structure and outputs a 128-dimensional feature vector corresponding to its final hidden state. This vector serves as a compact spatial–temporal representation, enriched with both appearance-based and motion-based information, and is used in subsequent stages such as feature fusion or classification.